OpenStudy (anonymous):

find the limit if it exists. i think i have to do something with a conjugate on this one. limg x->0 √2+x) -√2) / x

OpenStudy (anonymous):

gonna take a shot at it. be back in a minute

OpenStudy (klimenkov):

\[\lim_{x\rightarrow0}\frac{\sqrt{2+x}-\sqrt2}{x}\]

OpenStudy (anonymous):

yes just like that. still working on mine

hartnn (hartnn):

multiply and divide by \(\sqrt{2+x}+\sqrt2\)

hartnn (hartnn):

then use \((a-b)(a+b)=a^2-b^2\)

OpenStudy (anonymous):

i can't remember for the life of me how to multiply the numerators.

OpenStudy (anonymous):

im trying to use foil. by multiplying (√2x+x) -√2) (√2x+x)+√2), do i get rid of the radicals?

hartnn (hartnn):

as i already mentioned use :\((a-b)(a+b)=a^2-b^2\) here a= root(2+x), b =2

OpenStudy (klimenkov):

\[\lim_{x\rightarrow0}\frac{\sqrt{2+x}-\sqrt2}{x}=\lim_{x\rightarrow0}\frac{(\sqrt{2+x}-\sqrt2)(\sqrt{2+x}+\sqrt2)}{x(\sqrt{2+x}+\sqrt2)}=\lim_{x\rightarrow0}\frac{x}{x(\sqrt{2+x}+\sqrt2)}=\]\[\lim_{x\rightarrow0}\frac{1}{\sqrt{2+x}+\sqrt2}=\frac1{2\sqrt2}\]

OpenStudy (anonymous):

i got the same denom as klime

OpenStudy (klimenkov):

I'm sorry. I have problems with internet.

hartnn (hartnn):

yes,but did u get the numerator ?

OpenStudy (anonymous):

nope. that's where i was getting messed up

OpenStudy (anonymous):

gonna take a few to figure this out

OpenStudy (anonymous):

i have to get going :( i'll be back in half an hour! thanks for your help so far. gonna finish this when i get back